OpenStudy (zzr0ck3r):

let \(C\) be a closed convex subset of a Hilbert space \(\mathscr{H}\). Show that for each \(x\in\mathscr{H}\) there is a unique point \(y_0\in C\) such that \(||x-y_0||=\rho(x,C)=\inf\{||x-y|| \ : \ y\in C\}\)

OpenStudy (zzr0ck3r):

@eliassaab I am thinking we can have some sequence that converges to \(y_0\) since C is closed. s.t. we get the desired result. But as of now I have no idea how to go about it.

OpenStudy (zzr0ck3r):

Actually I think I found a good proof for this.

OpenStudy (zzr0ck3r):

ok the only other one that I have been having trouble with is A semi inner product on a linear space \(X\) is a function \(\langle ,\rangle :X \times X\rightarrow F\) satisfying the following conditions conditions a) \(\langle ax+by,z\rangle =a\langle x,z\rangle +b\langle y,z\rangle \) b) \(\langle x,y\rangle =\langle y,z\rangle \) if \(F=\mathbb{R}\) or \(\langle x,y\rangle =\overline{\langle y,x\rangle }\) if \(F=\mathbb{C}\). c) \(\langle x,x\rangle \ \ge \ 0\) d) \(\langle x,x\rangle =0 \) if \(x=0\) (note that non iff statement for this last property) I need to show that the following propositions hold, and I know that they hold for the case of an inner product. 1) \(\langle x+y, x+y\rangle =\langle x,x\rangle +2\mathscr{R}\langle x,y\rangle +\langle y,y\rangle \) 2) \(\lvert\langle x,y\rangle \rvert^2\ \le \ \langle x,x\rangle \langle y,y\rangle \)

OpenStudy (anonymous):

See this for 2) http://en.wikipedia.org/wiki/Cauchy%E2%80%93Schwarz_inequality

OpenStudy (anonymous):

\[ <x+y, x+y>=<x,x>+<x,y>+<y,x> +<y,y>=\\ <x,x>+<x,y>+\overline {<x,y>}+<y,y>= \langle x,x\rangle +2\mathscr{R}\langle x,y\rangle +\langle y,y\rangle \] Since \[ a+ i b + a-i b = 2a \]

OpenStudy (zzr0ck3r):

I think I might be figuring it out as we speak

OpenStudy (zzr0ck3r):

sweet:) tnx. I will keep trying without looking to hard at what you did, and know that I can come back to this if it gets to late:)

OpenStudy (zzr0ck3r):

@eliassaab that proof of the Cauchy inequality uses the non constrained definition of inner product, so they can claim that since \(v\ne0\) they can divide by \(\langle v, v \rangle \) but we do not have this property for the semi inner product

OpenStudy (zzr0ck3r):

sorry but what part of that? Its getting a little over my head.

OpenStudy (zzr0ck3r):

@eliassaab what about this one Let \(S\) be a subset of an inner product space \(X\). Show that \(S^{\perp}\) is a closed linear subspace of \(X\). I know how to show it is a linear subspace, that was easy enough, but what about the closed part?

OpenStudy (anonymous):

@zzr0ck3r . It is better to ask each question as a new problem, so new viewers would know which part we are talking about. Post it and tag me.